OpenStudy (anonymous):

Solve The Equation Algebraically 3x^2=12

OpenStudy (canadiansteve99):

what are the answer choices?

OpenStudy (anonymous):

+3 +2 3.5 1.5

OpenStudy (anonymous):

I got X=2

OpenStudy (anonymous):

How'd You Do It ?

OpenStudy (canadiansteve99):

Yeah x = 2. You multiply 3 and 2 and you get 6 right? Well then to equal 12, x has to equal 2, so the answer is x = 2.

OpenStudy (anonymous):

So On This On 2x^2=98 It Would Be 2*2 = 4 & Then To Equal 98 It Would Be x=94 ? @aaronq @CanadianSteve99

OpenStudy (canadiansteve99):

No, you would need to multiply 4 times a number to equal 98. If a variable (letter) is beside a number, like 2x = 18, then x would be 9 and you would multiply 2 and 9. You multiply 4 times x to equal 98.

OpenStudy (canadiansteve99):

Sorry if I over complicated that xD

OpenStudy (anonymous):

Yeah Just A Little @CanadianSteve99

OpenStudy (canadiansteve99):

Basically just multiply 4 times x, or divide 98 by 4.

OpenStudy (aaronq):

so when solving algebraic equations, you need to isolate the unknown variable (in this case it's x). you need to perform the inverse operations, for example if it's x/3=2, then you need to multiply both sides by 3. for you question: \(3x^2=12 \) divide both sides by 3: \(\dfrac{\cancel 3x^2}{\cancel 3}=\dfrac{12}{3} \) then take the square root of both sides: \(\sqrt x^2=\sqrt 4\) we're left with: \(x=2\)

OpenStudy (anonymous):

So The Square Root Of 12 ? @aaronq

OpenStudy (aaronq):

it's not the squared root of 12, but of \(\dfrac{12}{3}\) which is 4. so it's the squared root of 4, look at the first step

OpenStudy (anonymous):

Okay Okay So 12\ 3 = 4 & The Square Root Of 4 Is 2 Right ?

OpenStudy (aaronq):

yes and yes

OpenStudy (anonymous):

Okay I Think I Get It But Just To Make Sure If Its 2^2=98 Id Divide Both Sides By 2 Then Divide 98 & 2 I Got 49 & The Square Root Of 49 Is 7 Right ?

OpenStudy (aaronq):

was the original equation \(2x^2=98\) i think you missed the x in your expression.

OpenStudy (anonymous):

Yeah It Was

OpenStudy (aaronq):

than it is correct, good job on applying the process. \(\huge \checkmark\)

OpenStudy (aaronq):

then*

OpenStudy (anonymous):

Ahaa Thanks So Much , Couldnt Do It Without Your Help

OpenStudy (aaronq):

no problem, really! i gotta get going though, hopefully someone else can keep helping you, take care!

OpenStudy (anonymous):

Okay Thanks !

OpenStudy (anonymous):